OpenStudy (anonymous):

if y^2 - 3xy = x-5 find dy/dx

OpenStudy (bibby):

Implicit differentiation and product rule.

OpenStudy (anonymous):

?

OpenStudy (bibby):

Ok work through it step by step. What is the derivative of y^2?

OpenStudy (anonymous):

2y?

OpenStudy (bibby):

Yeah. If you learned the chain rule then you might remember that you then multiply by the derivative of the inside so it'd be\[2y \frac{ dy }{ dx? } = 2y*y'\] that's a y prime

OpenStudy (anonymous):

okay

OpenStudy (bibby):

Are you going to try the next part? \[\frac{d}{dx}3xy\]

OpenStudy (anonymous):

i don't understand how to do that...

OpenStudy (bibby):

Well remember that you can take constants out of the differentiation.\[\frac{d}{dx}3xy = 3\frac{d}{dx}xy\] and then use the product rule

OpenStudy (bibby):

back off bud this is mine

OpenStudy (shamil98):

y^2 - 3xy = x-5 Differentiate everything with respect to x, remember when you differentiate y though you end up it with it being multiplied by y prime (y'). For example you would make the derivative of y^2 into 2yy'. \[\huge y^2 - 3xy = x- 5\] To start off take the derivative of each value... Remember to use the product rule for 3xy

OpenStudy (shamil98):

Give it a try, and post your attempt .

OpenStudy (isaiah.feynman):

@bibby are you fighting to answer questions?

OpenStudy (bibby):

Nah. Me and shamil "shamil98" martinez have some playful scuffles over math. It's all in good fun, good nature, and good love

OpenStudy (anonymous):

i am still confused, i have never used this equation before

OpenStudy (anonymous):

\[y^2 - 3xy = x-5\] treat it like \[f^2(x) - 3xf(x) = x-5\]

OpenStudy (shamil98):

Take the derivative of every value separately \[\huge \frac{ d }{ dx } y^2 = 2yy'\] \[\huge \frac{ d }{ dx } -3xy = -3y -3xy'\] \[\huge \frac{ d }{ dx } x = 1\] \[\huge \frac{ d }{ dx } -5 = 0\] Put everything together. \[\huge 2yy' -3y - 3xy' = 1\] Group the y' and solve for y' \[\huge y'(2y-3x) = 3y+1\] \[\huge y' = \frac{ 3y+1 }{ 2y-3x }\]

OpenStudy (shamil98):

Do you understand how i got that?

OpenStudy (anonymous):

i think so i will try with the next one and see

OpenStudy (isaiah.feynman):

my boy @shamil98 delivers it again!

OpenStudy (anonymous):

sorry i still don't understand

OpenStudy (shamil98):

what don't you understand? ..

OpenStudy (bibby):

The multiplication rule goes something like u'v + uv'

OpenStudy (shamil98):

This gives a nice explanation of implicit differentiation, i found it useful. http://www.dummies.com/how-to/content/how-to-differentiate-implicitly0.html

OpenStudy (bibby):

Are you calling my lass a dummy

OpenStudy (shamil98):

no, the website just has good information with a poorly named domain.

OpenStudy (bibby):

That's what I thought

OpenStudy (shamil98):

adoptive fathers don't count

OpenStudy (anonymous):

but i don't understand what to do with xy shouldn't they be nonexistant because x^0 and y^0 don't exist?

OpenStudy (shamil98):

xy = x^1 * y^1 = xy.. when differentiating xy, you use the product rule: \[\huge f'g + fg'\]

OpenStudy (anonymous):

for y^2 +3x = xy -5 i got 3-y/x-2y

OpenStudy (shamil98):

\[\huge y^2 + 3x = xy -5\] \[\huge 2yy' + 3 = y + xy' \] \[\huge y'(2y-x) = y-3\] \[\huge y' = \frac{ y-3 }{ 2y-x }\]

OpenStudy (bibby):

\[2yy′+3=y+xy′\] \[yy′ -xy'=\frac{ y-3 }{ 2 }\] \[y'(y -x)=\frac{ y-3 }{ 2 }\] shouldn't it be 2y - 2x? or 2(y-x)

OpenStudy (shamil98):

2yy' + 3 = y+xy' 2yy' - xy' = y -3 y'(2y-x) = y-3 y' = y-3 / 2y - x you can't just take the 2 out -_-

OpenStudy (bibby):

shut your mouth. Don't give me a moronic smiley because you skipped a step of factoring. You've got this big fish in a little pond syndrome and it seriously gets under my skin brother

OpenStudy (shamil98):

Admit that you are wrong, young one. No shame in it , dear :)